### v1

# Tony Lindeberg

Abstract
This paper gives an in-depth theoretical analysis of the direction and speed selectivity properties of idealized models of the spatio-temporal receptive fields of simple cells and complex cells, based on the generalized Gaussian derivative model for visual receptive fields. According to this theory, the receptive fields are modelled as velocity-adapted affine Gaussian derivatives for different image velocities and different degrees of elongation.
By probing such idealized receptive field models of visual neurons to moving sine waves with different angular frequencies and image velocities, we characterize the computational models to a structurally similar probing method as is used for characterizing the direction and speed selective properties of biological neurons. It is shown that the direction selective properties become sharper with increasing order of spatial differentiation and increasing degree of elongation in the spatial components of the visual receptive fields. It is also shown that the speed selectivity properties are sharper for increasing order of spatial differentiation, while they are for the inclination angle $\theta=0$ independent of the degree of elongation.
By comparison to results of neurophysiological measurements of direction and speed selectivity for biological neurons in the primary visual cortex, we find that our theoretical results are qualitatively consistent with (i) velocity-tuned visual neurons that are sensitive to particular motion directions and speeds, and (ii) different visual neurons having broader vs. sharper direction and speed selective properties.
Our theoretical results in combination with results from neurophysiological characterizations of motion-sensitive visual neurons are also consistent with a previously formulated hypothesis that the simple cells in the primary visual cortex ought to be covariant under local Galilean transformations, so as to enable processing of visual stimuli with different motion directions and speeds.
journal: arXiv preprint
1 Introduction
In the quest of understanding the functional properties of the visual system, a key component concerns developing theoretical models for the computations performed in the visual pathways, and relating such computational models to neurophysiological measurements. Specifically, when formulating such theoretical models of visual processing, it is essential to understand the properties of these models and how those theoretical properties relate to corresponding properties as can be measured experimentally in biological vision.
Regarding the analysis of visual motion, it is in particular essential to understand the properties of the spatio-temporal receptive fields in the early visual pathway. The functional properties of the receptive fields of the visual neurons can aptly be regarded as the main primitives in computational modelling of the visual system. Reconstructing a full spatio-temporal receptive field does, however, constitute a non-trivial task, since it constitutes an inverse problem, which requires measurements of the response properties of each visual neuron to a large number of stimuli, and also a model of the receptive field to relate those measurements to. Probing the direction and speed selectivity properties of a visual neuron is, on the other hand, a much more straightforward task, which can be accomplished by moving a probing stimulus with different motion directions and different motion speeds within the support region of the receptive field. Consequently, there are substantially more available results concerning direction selectivity properties of visual neurons in the primary visual cortex (V1) and the middle temporal visual area (MT) (Hubel 1959, Hubel and Wiesel 1962, Emerson and Gerstein 1977, Albright 1984, Albright et al. 1984, Ganz and Felder 1984, Mikami et al. 1986a; 1986b, Orban et al. 1986, Lagae et al. 1993, Orban 1997, Livingstone 1998, Livingstone and Conway 2003, Born and Bradley 2005, Churchland et al. 2005, Gur et al. 2005, Moore et al. 2005, Priebe and Ferster 2005, Priebe et al. 2010, Wang and Yao 2011, Dai et al. 2025, Khamiss et al. 2025) than there are publicly available reconstructed spatio-temporal receptive fields (DeAngelis et al. 1995; 2004, de Valois et al. 2000). For this reason, it is of interest to establish theoretically based relationships between the properties of the spatio-temporal receptive fields of visual neurons and their direction selectivity properties, in order to bridge the gap between these two ways of characterizing motion-sensitive neurons.
The subject of this paper is to address this topic regarding the direction and speed selectivity properties of idealized mathematical models of spatio-temporal receptive fields in the primary visual cortex. For a theoretically well-founded way of modelling the spatio-temporal receptive fields of simple and complex cells for higher mammals, based on the generalized Gaussian derivative model for visual receptive fields (Lindeberg 2013; 2021), we will present an in-depth theoretical analysis of the direction and speed selectivity properties for idealized models of the spatio-temporal receptive fields of simple cells and complex cells based on this theory.
A main motivation for using this model for visual receptive fields, is that it is theoretically well motivated in the sense that the shapes of the corresponding spatio-temporal receptive fields can be derived by necessity, from structural properties of the environment in combination with internal consistency requirements to guarantee consistent treatment of image structures over multiple spatial and temporal scales.
A second major motivation is that the processing of spatio-temporal image data is integrated in the theory, which in this respect is different from the Gabor model for visual receptive fields, which has mainly been formulated for purely spatial image data. A third motivation is that the generalized Gaussian derivative theory for visual receptive field has the attractive property that it is able to handle the influence on image data due to geometric image transformations, with provable covariance properties under uniform scaling transformations, non-isotropic affine transformations, Galilean transformations and temporal scaling transformations (Lindeberg 2023b; 2025b; 2025d). Thereby, an idealized vision system based on receptive fields according to the generalized Gaussian derivative model for visual receptive fields can handle the variabilities in image data generated by varying the distance, the viewing direction and the relative motion between objects or spatio-temporal events for a visual agent that observes dynamic scenes. The receptive fields in this theory can also handle the variabilities caused by structurally similar spatio-temporal events occurring either either faster or slower relative to previously observed reference views.
Receptive fields generated from the generalized Gaussian derivative model for visual receptive fields have also been demonstrated to reasonably well model the qualitative shape of simple cells as recorded by DeAngelis et al. (1995; 2004), Conway and Livingstone (2006) and Johnson et al. (2008). See Figures 12–18 in Lindeberg (2021) for comparisons between biological receptive fields and idealized models thereof, based on the generalized Gaussian derivative model for visual receptive fields.
The methodology that we will follow in this treatment is to subject idealized models of simple and complex cells according to the generalized Gaussian derivative theory to similar probing mechanisms, as are used for probing the direction and speed selectivity properties of biological neurons. Specifically, we will focus on variabilities in the shapes of the spatio-temporal receptive fields, as induced by spanning the degrees of freedom of geometric image transformations, and analyze how these variabilities in the shapes of the spatio-temporal receptive fields affect the direction selectivity properties. In this way, we will provide an framework for how observed variabilities in the direction and speed selectivity properties of biological neurons could be explained in terms of variabilities in the shapes of the underlying simple cells, as well as the use of such simple cell primitives in our idealized models of complex cells.
This modelling will be performed for two types spatio-temporal receptive field models: (i) idealized models of simple cells in terms of non-separable velocity-adapted derivatives and (ii) idealized models of complex cells in terms of regionally integrated quasi quadrature combinations of non-separable velocity-adapted derivatives.
It will be shown that for the studied classes of idealized models of the spatio-temporal receptive fields, the direction selective properties depend strongly on the order of spatial differentiation and the degree of elongation of the spatial components of underlying receptive field models. It will also be shown that the speed selectivity properties depend strongly on the order of spatial differentiation, but not on the degree of elongation.
It will furthermore be shown that the results can be qualitatively related to neurophysiological results regarding the velocity sensitivity and direction selectivity properties of neurons in the primary visual cortex in monkeys, as well as to provide potential support for a previously formulated hypothesis that the spatio-temporal receptive fields of simple cells ought to be covariant under local Galilean transformations, so as to enable processing of visual stimuli with different motion directions and speeds.
In these ways, we do hence address the overreaching goal of bridging the gap between theoretical models and neurophysiological characterizations of spatio-temporal receptive fields in the primary visual cortex.
To promote opportunities for additional more detailed quantitative modelling of the spatio-temporal receptive fields in the primary visual cortex, we will also outline directions for further neurophysiological experiments.
1.1 Structure of this article
The presentation is organized as follows: After an overview of related work in Section 2, Section 3 begins by giving an overview of main components in the generalized Gaussian derivative theory of visual receptive fields for modelling the spatio-temporal receptive fields of simple cells and complex cells.
Section 4 then analyzes the direction and speed selectivity properties for 4 submodels of simple cells, based on velocity-adapted affine Gaussian derivatives of orders between 1 and 4, and corresponding to the order of spatial differentiation. Section 5 then analyzes the direction and speed selectivity properties for a set of different models of complex cells, in terms of regionally integrated quasi quadrature measures of the output from idealized models of simple cells up to order 4. A condensed summary of the derived results concerning direction and speed selectivity properties is given in Section 6.
Section 7 then compares the direction and speed selectivity properties that we have derived for our idealized models of simple cells and complex cells to corresponding results obtained from neurophysiological measurements in biological vision. Section 8 complements with a set of theoretically motivated questions for new neurophysiological experiments, to answer questions regarding more detailed computational modelling of the spatio-temporal receptive fields in the primary visual cortex. Finally, Section 9 concludes with a summary and discussion.
2 Relations to previous work
The direction selectivity properties of neurons in the primary visual cortex (V1) and the middle temporal visual area (MT) have been mapped in several neurophysiological studies by Hubel (1959), Hubel and Wiesel (1962), Emerson and Gerstein (1977), Albright 1984, Albright et al. (1984), Emerson and Gerstein (1977), Ganz and Felder (1984), Mikami et al. (1986a; 1986b), Orban et al. (1986), Lagae et al. (1993), Livingstone (1998), Livingstone and Conway (2003), Churchland et al. (2005), Gur et al. (2005), Moore et al. (2005), Priebe and Ferster (2005), Priebe et al. (2010), Wang and Yao (2011), Dai et al. (2025) and Khamiss et al. 2025) with excellent reviews by Orban (1997) and Born and Bradley (2005).
Computational models of motion analysis with close relations neurophysiological results regarding motion-sensitive neurons in the primary visual cortex and the middle temporal visual area have been formulated by Adelson and Bergen (1985), Heeger (1987), Wilson et al. (1992), Nowlan and Sejnowski (1995), Simoncelli and Heeger (1998) and Lisberger and Movshon (1999). A recent overview of experimentally driven computational models of area MT is given by Zarei Eskikand et al. (2024). The goal of the theoretical analysis in this paper is, however, not to address the motion estimation problem, but instead to understand how the direction and speed selectivity properties of visual neurons can be related to properties of the underlying spatio-temporal receptive fields, based on highly idealized models thereof.
Our knowledge about the functional properties of the receptive fields of simple cells in the primary visual cortex originates from the pioneering work by Hubel and Wiesel (1959; 1962; 1968; 2005) followed by more detailed characterizations by DeAngelis et al. (1995; 2004), de Valois et al. (2000), Ringach (2002; 2004), Conway and Livingstone (2006), Johnson et al. (2008), Walker et al. (2019) and De and Horwitz (2021).
Computational models of simple cells have specifically been expressed in terms of Gabor filters by Marcelja (1980), Jones and Palmer (1987a; 1987b), Porat and Zeevi (1988), Ringach (2002; 2004), Serre et al. (2007), Baspinar et al. (2018; 2020) and De and Horwitz (2021), and in terms of Gaussian derivatives by Koenderink and van Doorn (1984; 1987; 1992), Young (1987), Young et al. (2001; 2001) and Lindeberg (2013; 2021). Theoretical models of early visual processes have also been formulated based on Gaussian derivatives by Lowe (2000), May and Georgeson (2007), Hesse and Georgeson (2005), Georgeson et al. (2007), Hansen and Neumann (2008), Wallis and Georgeson (2009), Wang and Spratling (2016), Pei et al. (2016), Ghodrati et al. (2017), Kristensen and Sandberg (2021), Abballe and Asari (2022), Ruslim et al. (2023) and Wendt and Faul (2024). Since Gabor models have been primarily applied to a purely spatial domain, and not extensively studied regarding spatio-temporal data, we will fully focus on the spatio-temporal receptive fields according to the generalized Gaussian derivative model in this treatment.
The neurophysiological properties of complex cells have been studied by Movshon et al. (1978), Emerson et al. (1987), Martinez and Alonso (2001), Touryan et al. (2002; 2005), Rust et al. (2005), van Kleef et al. (2010), Goris et al. (2015), Li et al. (2015) and Almasi et al. (2020), as well as been modelled computationally by Adelson and Bergen (1985), Heeger (1992), Serre and Riesenhuber (2004), Einhäuser et al. (2002), Kording et al. (2004), Merolla and Boahen (2004), Berkes and Wiscott (2005), Carandini (2006), Hansard and Horaud (2011), Franciosini et al. (2019), Lindeberg (2020), Lian et al. (2021), Oleskiw et al. (2024), Yedjour and Yedjour (2024) and Nguyen et al. (2024).
Notably, in relation to the generalized quadratic models of complex cells in V1 to be considered in this paper, Rowekamp and Sharpee (2025) have found that quadratic computations strongly increase both the predictive power of their models of visual neurons in V1, V2 and V4 as well as their neural selectivity to natural stimuli.
More explicit neural models of spatio-temporal receptive fields in the primary visual cortex have also been formulated by Heitmann and Ermentrout (2016), Chizhov and Merkulyeva (2020), Chariker et al. (2021; 2022), Freeman (2021) and Larisch and Hamker (2025), including studies of direction selectivity properties.
Compared to such explicit neural models of spatio-temporal receptive fields, a conceptual difference with our approach is additionally that we instead analyze compact functional models of the spatio-temporal receptive fields, which have been determined from an axiomatically determined normative theory of visual receptive fields. In this way, we are aiming at more compact descriptions of the functional phenomena, as parameterized by a small set of shape parameters of the spatio-temporal receptive fields, in contrast to instead having to determine a much larger number of filter weights and other dynamics parameters in the more explicit neural models and computing the results by numerical simulations.
The orientation selectivity properties of idealized models of simple and complex cells have been analyzed in detail in Lindeberg (2025a; 2025c). The notions of direction selectivity and speed selectivity, that we address in this work, are related to the notion of orientation selectivity, but are structurally different properties. The notion of orientation selectivity concerns behaviour over a spatial domain, whereas the notions of direction and speed selectivity concern behaviour over the joint spatio-temporal domain.
In this work, we theoretically analyze the results of how the output from idealized models of simple cells and complex cells are affected by varying both the direction and the speed of motion stimuli. We also relate these results to direction and speed selectivity properties of biological neurons, and to a hypothesis that the receptive fields in the primary visual cortex ought to be covariant under local Galilean transformations.
The latter will then imply a variability of the shapes of the spatio-temporal receptive fields under the degrees of freedom of local Galilean transformations. In practice, this would imply that the there would be multiple copies of similar types of spatio-temporal receptive fields for different values of the velocity parameter $v$ in the idealized model (3.1) of simple cells, as illustrated in Figures 2 and 2.
3 The generalized Gaussian derivative model for visual receptive fields
In this section, we will describe the idealized models of simple cells and complex cells, that we will then analyze the direction selectivity properties for in Sections 4 and 5.
v=1v=-1 v=1/2v=-1/2 $v=0$ $v=1/2$ $v=1$
$\scriptsize{\sigma_{x}=2,\sigma_{t}=2}$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
$\scriptsize{\sigma_{x}=1,\sigma_{t}=2}$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
$\scriptsize{\sigma_{x}=2,\sigma_{t}=1}$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
$\scriptsize{\sigma_{x}=1,\sigma_{t}=1}$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 1: Non-causal joint spatio-temporal receptive fields over a 1+1D spatio-temporal domain in terms of the second-order spatial derivative of the form $T_{xx,\text{norm}}(x,t;\;s,\tau,v)$ according to (3.1) of the product of a velocity-adapted 1-D Gaussian kernel over the spatial domain and the non-causal temporal kernel over the temporal domain according to (3). The spatio-temporal receptive fields are shown for different values of the spatial scale parameter $\sigma_{x}=\sqrt{s}$ and the temporal scale parameter $\sigma_{t}=\sqrt{\tau}$ in dimensions of $\mbox{length}]$ and $\mbox{time}]$. (Horizontal axes: Spatial image coordinate $x\in[-10,10]$. Vertical axes: Temporal variable $t\in[-4,4]$.)
v=1v=-1 v=1/2v=-1/2 $v=0$ $v=1/2$ $v=1$
$\scriptsize{\sigma_{x}=2,\sigma_{t}=2}$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
$\scriptsize{\sigma_{x}=1,\sigma_{t}=2}$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
$\scriptsize{\sigma_{x}=2,\sigma_{t}=1}$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
$\scriptsize{\sigma_{x}=1,\sigma_{t}=1}$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 2: Time-causal joint spatio-temporal receptive fields over a 1+1D spatio-temporal domain in terms of the second-order spatial derivative of the form $T_{xx,\text{norm}}(x,t;\;s,\tau,v)$ according to (3.1) of the product of a velocity-adapted 1-D Gaussian kernel over the spatial domain and the time-causal limit kernel over the temporal domain according to (4). The spatio-temporal receptive fields are shown for different values of the spatial scale parameter $\sigma_{x}=\sqrt{s}$ and the temporal scale parameter $\sigma_{t}=\sqrt{\tau}$ in dimensions of $\mbox{length}]$ and $\mbox{time}]$. (Horizontal axes: Spatial image coordinate $x\in[-10,10]$. Vertical axes: Temporal variable $t\in[0,8]$.)
3.1 Idealized model for simple cells
According to the generalized Gaussian derivative model for visual receptive fields, linear spatio-temporal receptive fields corresponding to the simple cells in the primary visual cortex are modelled by spatio-temporal derivatives of velocity-adapted affine Gaussian kernels in the following idealized way (Lindeberg 2011; 2013; 2016; 2021):
$\begin{split}T_{\text{simple}}(x_{1},x_{2},t;\;\sigma_{\varphi},\sigma_{t},\varphi,v,\Sigma_{\varphi},m,n)\end{split}$
$\begin{split}&=T_{{\varphi}^{m},{\bar{t}}^{n},\scriptsize\mbox{norm}}(x_{1},x_{2},t;\;\sigma_{\varphi},\sigma_{t},v,\Sigma_{\varphi})\end{split}$
$\begin{split}&=\sigma_{\varphi}^{m}\,\sigma_{t}^{n}\,\partial_{\varphi}^{m}\,\partial_{\bar{t}}^{n}\left(g(x_{1}-v_{1}t,x_{2}-v_{2}t;\;\Sigma_{\varphi})\,h(t;\;\sigma_{t})\right),\end{split}$ (1)
where
  • $\varphi\in[-\pi,\pi]$ is the preferred orientation of the receptive field,
  • $\sigma_{\varphi}\in{\mathbb{R}}_{+}$ is the amount of spatial smoothing (in units of the spatial standard deviation),
  • $\partial_{\varphi}^{m}=(\cos\varphi\,\partial_{x_{1}}+\sin\varphi\,\partial_{x_{2}})^{m}$ is an $m$:th-order directional derivative operator in the direction $\varphi$,
  • $\Sigma_{\varphi}$ is a $2\times 2$ symmetric positive definite covariance matrix, with one of its eigenvectors in the direction of $\varphi$,
  • $g(x;\;\Sigma_{\varphi})$ is a 2-D affine Gaussian kernel, with its shape determined by the spatial covariance matrix $\Sigma_{\varphi}$
    $g(x;\;\Sigma_{\varphi})=\frac{1}{2\pi\sqrt{\det\Sigma_{\varphi}}}e^{-x^{T}\Sigma_{\varphi}^{-1}x/2}$ (2)
    for $x=(x_{1},x_{2})^{T}\in{\mathbb{R}}^{2}$.
  • $\sigma_{t}$ represents the amount of temporal smoothing (in units of the temporal standard deviation),
  • $v=(v_{1},v_{2})^{T}$ represents a local motion vector in the direction $\varphi$ of the spatial orientation of the receptive field,
  • $\partial_{\bar{t}}^{n}=(\partial_{t}+v_{1}\,\partial_{x_{1}}+v_{2}\,\partial_{x_{2}})^{n}$ represents an $n$:th-order velocity-adapted temporal derivative operator, and
  • $h(t;\;\sigma_{t})$ represents a temporal smoothing kernel with temporal standard deviation $\sigma_{t}$.
In the case when the temporal domain is regarded as non-causal (implying that the future relative to any temporal moment can be accessed, as it can be for pre-recorded video data), the temporal kernel can be chosen as the 1-D Gaussian kernel
$h(t;\;\sigma_{t})=\frac{1}{\sqrt{2\pi}\sigma_{t}}\,e^{-t^{2}/2\sigma_{t}^{2}},$ (3)
whereas in the case when the temporal is truly time-causal (corresponding to the more realistic real-time scenario, where the future cannot be accessed), the temporal kernel can determined as the time-causal limit kernel (Lindeberg 2016 Section 5; Lindeberg 2023a Section 3)
$h(t;\;\sigma_{t})=\psi(t;\;\sigma_{t},c),$ (4)
characterized by having a Fourier transform of the form
$\hat{\Psi}(\omega;\;\sigma_{t},c)=\prod_{k=1}^{\infty}\frac{1}{1+i\,c^{-k}\sqrt{c^{2}-1}\,\sigma_{t}\,\omega}.$ (5)
Figure 2 shows examples of such receptive fields over a non-causal 1+1-D spatio-temporal domain, for the case when the spatial differentiation order is $m=2$ and the temporal differentiation order is $n=0$. Figure 2 shows corresponding examples of such receptive fields over a time-causal 1+1-D spatio-temporal domain, for the same orders of spatial and temporal differentiation.
3.2 Idealized models for complex cells based on velocity-tuned simple cells
When to define idealized models of complex cells from the output from simple cells, one basic fact to consider is that if one applies velocity-adapted spatio-temporal receptive fields for non-zero orders $n$ of temporal differentiation, then the resulting temporal derivatives will be zero if the velocity parameter $v$ of the spatio-temporal receptive field is equal to the image velocity $u$ of the image stimulus. For this reason, it we want to have a model of complex cells, that is to respond to image data with large values for matching image velocities, without introducing a magnitude inversion step, it is natural to formulate an idealized model of complex cells based on the output from simple cells for zero order $n=0$ of temporal differentiation.
Based on Lindeberg (2025a) Equation (23), with the complementary scale normalization parameter $\Gamma$ in that more general expression set to $\Gamma=0$, for simplicity, we will consider the following quasi quadrature combination of first- and second-order spatial derivatives
$({\cal Q}_{\varphi,12,\scriptsize\mbox{vel},\scriptsize\mbox{norm}}L)^{2}=L_{\varphi,\scriptsize\mbox{norm}}^{2}+\,C_{\varphi}\,L_{\varphi\varphi,\scriptsize\mbox{norm}}^{2}$ (6)
and the following new quasi quadrature combination of third- and fourth-order spatial derivatives
$({\cal Q}_{\varphi,34,\scriptsize\mbox{vel},\scriptsize\mbox{norm}}L)^{2}=L_{\varphi\varphi\varphi,\scriptsize\mbox{norm}}^{2}+\,C_{\varphi}\,L_{\varphi\varphi\varphi\varphi,\scriptsize\mbox{norm}}^{2},$ (7)
where the individual components in this expression are defined from velocity-adapted spatio-temporal receptive fields according to
$L_{\varphi,\scriptsize\mbox{norm}}(\cdot,\cdot,\cdot;\;\sigma_{\varphi},\sigma_{t},v,\Sigma_{\varphi})=\\ =T_{\varphi,\scriptsize\mbox{norm}}(\cdot,\cdot,\cdot;\;\sigma_{\varphi},\sigma_{t},v,\Sigma_{\varphi})*f(\cdot,\cdot,\cdot),$ (8)
$L_{\varphi\varphi,\scriptsize\mbox{norm}}(\cdot,\cdot,\cdot;\;\sigma_{\varphi},\sigma_{t},v,\Sigma_{\varphi})=\\ =T_{\varphi\varphi,\scriptsize\mbox{norm}}(\cdot,\cdot,\cdot;\;\sigma_{\varphi},\sigma_{t},v,\Sigma_{\varphi})*f(\cdot,\cdot,\cdot),$ (9)
and
$L_{\varphi\varphi\varphi,\scriptsize\mbox{norm}}(\cdot,\cdot,\cdot;\;\sigma_{\varphi},\sigma_{t},v,\Sigma_{\varphi})=\\ =T_{\varphi\varphi\varphi,\scriptsize\mbox{norm}}(\cdot,\cdot,\cdot;\;\sigma_{\varphi},\sigma_{t},v,\Sigma_{\varphi})*f(\cdot,\cdot,\cdot),$ (10)
$L_{\varphi\varphi\varphi\varphi,\scriptsize\mbox{norm}}(\cdot,\cdot,\cdot;\;\sigma_{\varphi},\sigma_{t},v,\Sigma_{\varphi})=\\ =T_{\varphi\varphi\varphi\varphi,\scriptsize\mbox{norm}}(\cdot,\cdot,\cdot;\;\sigma_{\varphi},\sigma_{t},v,\Sigma_{\varphi})*f(\cdot,\cdot,\cdot),$ (11)
with the underlying velocity-adapted spatio-temporal receptive fields $T_{\varphi^{m},t^{n},\scriptsize\mbox{norm}}(x_{1},x_{2},t;\;\sigma_{\varphi},\sigma_{t},v,\Sigma_{\varphi})$ according to (3.1) for $n=0$.
While the quasi quadrature measures ${\cal Q}_{\varphi,12,\scriptsize\mbox{vel},\scriptsize\mbox{norm}}L$ and ${\cal Q}_{\varphi,34,\scriptsize\mbox{vel},\scriptsize\mbox{norm}}L$ have been designed to reduce the spatial variations in the underlying pairs of first-order and second-order or third-order and fourth-order spatial derivative responses, as approximations to the notion of quadrature pairs, we will here instead consider corresponding extensions to spatially integrated quasi quadrature measures according to
$((\overline{{\cal Q}}_{\varphi,12,\scriptsize\mbox{vel},\scriptsize\mbox{norm}}L)(\cdot,\cdot,t;\;\sigma_{\varphi},\sigma_{t},v,\Sigma_{\varphi}))^{2}=\\ =g(\cdot,\cdot;\;\gamma^{2}\Sigma_{\varphi})*(({\cal Q}_{\varphi,12,\scriptsize\mbox{vel},\scriptsize\mbox{norm}}L)(\cdot,\cdot,t;\;\sigma_{\varphi},\sigma_{t},v,\Sigma_{\varphi}))^{2}$ (12)
and
$((\overline{{\cal Q}}_{\varphi,34,\scriptsize\mbox{vel},\scriptsize\mbox{norm}}L)(\cdot,\cdot,t;\;\sigma_{\varphi},\sigma_{t},v,\Sigma_{\varphi}))^{2}=\\ =g(\cdot,\cdot;\;\gamma^{2}\Sigma_{\varphi})*(({\cal Q}_{\varphi,34,\scriptsize\mbox{vel},\scriptsize\mbox{norm}}L)(\cdot,\cdot,t;\;\sigma_{\varphi},\sigma_{t},v,\Sigma_{\varphi}))^{2},$ (13)
where $\gamma>0$ is a relative integration scale, that we will henceforth choose as $\gamma=\sqrt{2}$. In this way, the results will have a much lower variability with respect to the spatial positions in the image domain.
The underlying motivation for these constructions, is that the resulting spatio-temporal quasi quadrature measures
$\overline{{\cal Q}}_{\varphi,12,\scriptsize\mbox{vel},\scriptsize\mbox{norm}}L$ and $\overline{{\cal Q}}_{\varphi,34,\scriptsize\mbox{vel},\scriptsize\mbox{norm}}L$ should then assume their maximum values when the velocity parameter $v$ of the receptive field is equal to the image velocity $u$ of the input stimulus.
4 Direction and speed selective properties for idealized models of simple cells
For probing the direction selectivity properties of the different forms of spatio-temporal receptive fields described in Section 3, we will in this paper throughout make use of stimuli in terms of moving sine waves of the form
$f(x_{1},x_{2},t)=\\ =\sin\left(\omega\cos(\theta)\,x_{1}+\omega\sin(\theta)\,x_{2}-u\,t+\beta\right),$ (14)
where $\omega$ is the magnitude of the angular frequency, $\theta$ its inclination angle, $u$ is the image velocity, and $\beta$ is a phase angle, see Figure 3 for an illustration.
Figure 3: Schematic illustration of the modelling situation studied in the theoretical analysis, where the coordinate system is aligned to the preferred orientation $\varphi=0$ of the receptive field, and the receptive field is then exposed to a moving sine wave pattern with inclination angle $\theta$. In this figure, the sine wave pattern is schematically illustrated by a set of level lines, overlayed onto a few level curves of a first-order affine Gaussian derivative kernel. (Horizontal axis: spatial coordinate $x_{1}$. Vertical axis: spatial coordinate $x_{2}$.)
While it may be more common in some neurophysiological experiments to instead probe the direction selectivity properties of visual neurons with moving bars, a main theoretical reason for using moving sine waves is that the responses of the spatio-temporal receptive fields can then be computed in closed form, for the spatio-temporal receptive fields based on the generalized Gaussian derivative model.
Furthermore, to make the results independent of the relationship between the spatial and temporal scale levels used in the models of the spatio-temporal receptive fields, we will
  • for each receptive field choose the angular frequency of the probing sine wave in such a way that it maximizes the response of the spatio-temporal receptive field, and
  • for the spatio-temporal receptive fields with velocity adaptation, show the direction and speed selectivity graphs only for the special case when the spatial and the temporal scale levels are coupled in relation to the magnitude $v$ of the velocity parameter $v$ of the receptive field according to
    $\frac{\sigma_{x}}{\sigma_{t}}=v,$ (15)
    which is natural from the viewpoint of regarding the velocity estimation problem for a visual neuron as depending on the extent of the underlying receptive field over the spatial and the temporal domains.
In these ways, the resulting measures of the direction selectivity properties of the idealized models of the spatio-temporal receptive fields can be regarded as reflecting inherent properties of their computational functions.
To handle the possible reservation, in case the coupling of the spatial and temporal scale parameters $\sigma_{1}$ and $\sigma_{t}$ according to (15) would constitute a too strong simplification, we do, however, provide explicit mathematical models for the combined direction and speed selectivity properties for the idealized models of simple cells in Section 6.
Concerning the types of idealized models of motion-sensitive neurons, we will perform separate analyses of simple cells for different orders $m$ of spatial differentiation in this section. Then, in Section 5, we will perform corresponding analyses of direction and speed selectivity properties for a set of models of complex cells. The motivation for studying such a large variety of models of visual neurons is that in a neurophysiological experiment, where the direction and speed selectivity properties of a neuron in the primary visual cortex, it may not be a priori clear if that neuron would correspond to a simple cells or a complex cell, and furthermore regarding the models of simple cells, it may not be a priori clear what order of spatial differentiation the neuron would correspond to. Thereby, it is necessary to analyze a rather large set of idealized models of simple cells and complex cells with a similar theoretical probing scheme.
The reader more interested in the results than the details of the derivations can proceed directly to Section 6, where a condensed summary is given of some of the main results.
4.1 Analysis for first-order simple cell without temporal differentiation
Consider a velocity-adapted receptive field corresponding to a first-order scale-normalized Gaussian derivative with scale parameter $\sigma_{1}$ and velocity $v$ in the horizontal $x_{1}$-direction, a zero-order Gaussian kernel with scale parameter $\sigma_{2}$ in the vertical $x_{2}$-direction, and a zero-order Gaussian derivative with scale parameter $\sigma_{t}$ in the temporal direction, corresponding to $\varphi=0$, $v=0$, $\Sigma_{0}=\operatorname{diag}(\sigma_{1}^{2},\sigma_{2}^{2})$, $m=1$ and $n=0$ in (3.1):
$\begin{split}&T_{0,\scriptsize\mbox{norm}}(x_{1},x_{2},t;\;\sigma_{1},\sigma_{2},\sigma_{t})=\end{split}$
$\begin{split}&=\frac{\sigma_{1}}{(2\pi)^{3/2}\,\sigma_{1}\sigma_{2}\sigma_{t}}\,\partial_{x_{1}}\left.\left(e^{-x_{1}^{2}/2\sigma_{1}^{2}-x_{2}^{2}/2\sigma_{2}^{2}-t^{2}/2\sigma_{t}^{2}}\right)\right|_{x_{1}\rightarrow x_{1}-vt}\end{split}$
$\begin{split}&=\frac{(x_{1}-vt)}{(2\pi)^{3/2}\,\sigma_{1}^{2}\sigma_{2}\sigma_{t}}\,e^{-(x_{1}-vt)^{2}/2\sigma_{1}^{2}-x_{2}^{2}/2\sigma_{2}^{2}-t^{2}/2\sigma_{t}^{2}}.\end{split}$ (16)
The corresponding receptive field response is then, after solving the convolution integral in Mathematica,
$\begin{split}L_{0,\scriptsize\mbox{norm}}(x_{1},x_{2},t;\;\sigma_{1},\sigma_{2},\sigma_{t})=\end{split}$
$\begin{split}&=\int_{\xi_{1}=-\infty}^{\infty}\int_{\xi_{2}=-\infty}^{\infty}\int_{\zeta=-\infty}^{\infty}T_{0,\scriptsize\mbox{norm}}(\xi_{1},\xi_{2},\zeta;\;\sigma_{1},\sigma_{2},\sigma_{t})\end{split}$
[…]
$\begin{split}&=\omega\,\sigma_{1}\cos\theta\,\end{split}$
$\begin{split}&\phantom{==}\times e^{-\frac{\omega^{2}}{2}\left(\left(\sigma_{1}^{2}+\sigma_{t}^{2}v^{2}\right)\cos^{2}(\theta)+\sigma_{2}^{2}\sin^{2}\theta-2\sigma_{t}^{2}uv\cos\theta+\sigma_{t}^{2}u^{2}\right)}\end{split}$
$\begin{split}&\phantom{==}\times\cos\left(\cos(\theta)\,x_{1}+\sin(\theta)\,x_{2}-\omega\,u\,t+\beta\right),\end{split}$ (17)
i.e., a cosine wave with amplitude
$\begin{split}&A_{\varphi}(\theta,u,\omega;\;\sigma_{1},\sigma_{2},\sigma_{t},v)=\end{split}$
$\begin{split}&=\omega\,\sigma_{1}\left|\cos\theta\right|\,\end{split}$
$\begin{split}&\phantom{==}\times e^{-\frac{\omega^{2}}{2}\left(\cos^{2}\theta\left(\sigma_{1}^{2}+\sigma_{t}^{2}v^{2}\right)+\sigma_{2}^{2}\sin^{2}\theta-2\sigma_{t}^{2}uv\cos\theta+\sigma_{t}^{2}u^{2}\right)}.\end{split}$ (18)
Assume that a biological experiment regarding the response properties of the receptive field is performed by varying the angular frequency $\omega$ to get the maximum value of the response over these parameters. Differentiating the amplitude $A_{\varphi}$ with respect to $\omega$ and setting this derivative to zero, while also setting $\sigma_{2}=\kappa\,\sigma_{1}$ and $v=\sigma_{1}/\sigma_{t}$, then gives
$\hat{\omega}_{\varphi}=\frac{1}{\sqrt{\frac{\sigma_{1}^{2}\left(\kappa^{2}v^{2}\sin^{2}(\theta)-2uv\cos(\theta)+2v^{2}\cos^{2}(\theta)+u^{2}\right)}{v^{2}}}}$ (19)
Inserting this value into $A_{\varphi}(\theta,u,\omega;\;\sigma_{1},\sigma_{2},\sigma_{t},v)$, then gives the following combined direction and speed selectivity measure
$A_{\varphi,\max}(\theta,u;\;\kappa,v)=\\ =\frac{\cos(\theta)}{\sqrt{e}\sqrt{\kappa^{2}\sin^{2}(\theta)-\frac{2u\cos(\theta)}{v}+2\cos^{2}(\theta)+\frac{u^{2}}{v^{2}}}},$ (20)
which for the inclination angle $\theta=0$ and for a coupling of the speed values of the form u=rvu=r\,v, reduces to the following speed dependency, notably independent of the degree of elongation $\kappa$:
$R_{\varphi}(r)=A_{\varphi,\max}(0,r\,v;\;\kappa,r\,v)=\frac{1}{\sqrt{r^{2}-2r+2}}.$ (21)
First-order simple cell Second-order simple cell Third-order simple cell Fourth-order simple cell
$\kappa=1$ Refer to caption Refer to caption Refer to caption Refer to caption
$\kappa=2$ Refer to caption Refer to caption Refer to caption Refer to caption
$\kappa=4$ Refer to caption Refer to caption Refer to caption Refer to caption
$\kappa=8$ Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: Graphs of the direction selectivity for velocity-tuned models of simple cells based on (left column) first-order directional derivatives of affine Gaussian kernels combined with zero-order temporal Gaussian kernels, (middle left column) second-order directional derivatives of affine Gaussian kernels combined with zero-order temporal Gaussian kernels, (middle right column) third-order directional derivatives of affine Gaussian kernels combined with zero-order temporal Gaussian kernels, or (right column) fourth-order directional derivatives of affine Gaussian kernels combined with zero-order temporal Gaussian kernels, shown for different values of the degree of elongation $\kappa$ between the spatial scale parameters in the vertical vs. the horizontal directions. Observe how the direction selectivity varies strongly depending on the eccentricity $\epsilon=1/\kappa$ of the receptive fields. (top row) Results for $\kappa=1$. (second row) Results for $\kappa=2$. (third row) Results for $\kappa=4$. (bottom row) Results for $\kappa=8$.
First-order simple cell Second-order simple cell Third-order simple cell Fourth-order simple cell
$r=1/$ Refer to caption Refer to caption Refer to caption Refer to caption
$r=1/2$ Refer to caption Refer to caption Refer to caption Refer to caption
$r=1$ Refer to caption Refer to caption Refer to caption Refer to caption
$r=2$ Refer to caption Refer to caption Refer to caption Refer to caption
$r=4$ Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: Graphs of the direction selectivity for velocity-tuned models of simple cells based on (left column) first-order directional derivatives of affine Gaussian kernels combined with zero-order temporal Gaussian kernels, (middle left column) second-order directional derivatives of affine Gaussian kernels combined with zero-order temporal Gaussian kernels, (middle right column) third-order directional derivatives of affine Gaussian kernels combined with zero-order temporal Gaussian kernels, or (right column) fourth-order directional derivatives of affine Gaussian kernels combined with zero-order temporal Gaussian kernels, shown for different values of the ratio $r$ in the relationship u=rvu=r\,v between the speed $u$ of the motion stimulus and the speed $v$ of the receptive field, for a fixed value of $\kappa=2$ between the spatial scale parameters in the vertical vs. the horizontal directions. Observe how the direction selectivity varies strongly depending on the eccentricity $\epsilon=1/\kappa$ of the receptive fields. (top row) Results for $r=1/4$. (second row) Results for $r=1/2$. (third row) Results for $r=1$. (fourth row) Results for $r=2$. (bottom row) Results for $r=4$.
First-order simple cell Second-order simple cell Third-order simple cell Fourth-order simple cell
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 6: Graphs of the speed selectivity curves $R_{\varphi^{m}}(r)$ according to (21), (27), (33) and (39) for the inclination angle $\theta=0$ for velocity-tuned models of simple cells based on (left) first-order directional derivatives of affine Gaussian kernels combined with zero-order temporal Gaussian kernels, (middle left) second-order directional derivatives of affine Gaussian kernels combined with zero-order temporal Gaussian kernels, (middle right) third-order directional derivatives of affine Gaussian kernels combined with zero-order temporal Gaussian kernels, or (right) fourth-order directional derivatives of affine Gaussian kernels combined with zero-order temporal Gaussian kernels, shown for different values of the parameter $r$ in the relationship u=rvu=r\,v between the speed $u$ of the motion stimulus and the speed $v$ of the receptive field. Notably, these measures are independent of the degree of elongation $\kappa$ of the spatio-temporal receptive fields. (Horizontal axes: relative velocity parameter $r\in[0,4]$. Vertical axes: velocity sensitivity function $R_{\varphi^{m}}(r)$.)
The left column in Figure 4 shows the result of plotting the measure $A_{\varphi,\max}(\theta,u;\;\kappa,v)$ of the direction selectivity as function of the inclination angle $\theta$ for a few values of the scale parameter ratio $\kappa$ for the special case when the speed of the stimulus is equal to the speed of the velocity $u=v$, with the values rescaled such that the peak value for each graph is equal to 1. As we can see from the graphs, the direction selectivity becomes sharper with increasing values of the degree of elongation $\kappa$. The left column in Figure 5 shows corresponding results of varying the magnitude of the motion according to u=rvu=r\,v for different values of $r$, while keeping the degree of elongation fixed to $\kappa=2$. As we can see from the results, the direction selectivity of the simple cells prefers motion speeds $u$ that are near the preferred speed $v$ of the receptive field. The left graph in Figure 6 finally shows the relative velocity sensitivity curve $R_{\varphi}(r)$.
4.2 Analysis for second-order simple cell without temporal differentiation
Consider next a velocity-adapted receptive field corresponding to a second-order scale-normalized Gaussian derivative with scale parameter $\sigma_{1}$ and velocity $v$ in the horizontal $x_{1}$-direction, a zero-order Gaussian kernel with scale parameter $\sigma_{2}$ in the vertical $x_{2}$-direction, and a zero-order Gaussian derivative with scale parameter $\sigma_{t}$ in the temporal direction, corresponding to $\varphi=0$, $v=0$, $\Sigma_{0}=\operatorname{diag}(\sigma_{1}^{2},\sigma_{2}^{2})$, $m=2$ and $n=0$ in (3.1):
$\begin{split}&T_{00,\scriptsize\mbox{norm}}(x_{1},x_{2},t;\;\sigma_{1},\sigma_{2},\sigma_{t})=\end{split}$
$\begin{split}&=\frac{\sigma_{1}^{2}}{(2\pi)^{3/2}\,\sigma_{1}\sigma_{2}\sigma_{t}}\,\partial_{x_{1}x_{1}}\left.\left(e^{-x_{1}^{2}/2\sigma_{1}^{2}-x_{2}^{2}/2\sigma_{2}^{2}-t^{2}/2\sigma_{t}^{2}}\right)\right|_{x_{1}\rightarrow x_{1}-vt}\end{split}$
$\begin{split}&=\frac{((x_{1}-vt)^{2}-\sigma_{1}^{2})}{(2\pi)^{3/2}\,\sigma_{1}^{3}\sigma_{2}\sigma_{t}}\,e^{-(x_{1}-vt)^{2}/2\sigma_{1}^{2}-x_{2}^{2}/2\sigma_{2}^{2}-t^{2}/2\sigma_{t}^{2}}.\end{split}$ (22)
The corresponding receptive field response is then, after solving the convolution integral in Mathematica,
$\begin{split}L_{00,\scriptsize\mbox{norm}}(x_{1},x_{2},t;\;\sigma_{1},\sigma_{2},\sigma_{t})=\end{split}$
$\begin{split}&=\int_{\xi_{1}=-\infty}^{\infty}\int_{\xi_{2}=-\infty}^{\infty}\int_{\zeta=-\infty}^{\infty}T_{00,\scriptsize\mbox{norm}}(\xi_{1},\xi_{2},\zeta;\;\sigma_{1},\sigma_{2},\sigma_{t})\end{split}$
[…]
$\begin{split}&=-\omega^{2}\sigma_{1}^{2}\cos^{2}\theta\,\end{split}$
$\begin{split}&\phantom{==}\times e^{-\frac{\omega^{2}}{2}\left(\left(\sigma_{1}^{2}+\sigma_{t}^{2}v^{2}\right)\cos^{2}\theta+\sigma_{2}^{2}\sin^{2}\theta-2\sigma_{t}^{2}uv\cos\theta+\sigma_{t}^{2}u^{2}\right)}\end{split}$
$\begin{split}&\phantom{==}\times\sin\left(\cos(\theta)\,x_{1}+\sin(\theta)\,x_{2}-\omega\,u\,t+\beta\right),\end{split}$ (23)
i.e., a sine wave with amplitude
[…]
$\begin{split}&=\omega^{2}\sigma_{1}^{2}\cos^{2}\theta\,\end{split}$
$\begin{split}&\phantom{==}\times e^{-\frac{\omega^{2}}{2}\left(\cos^{2}\theta\left(\sigma_{1}^{2}+\sigma_{t}^{2}v^{2}\right)+\sigma_{2}^{2}\sin^{2}\theta-2\sigma_{t}^{2}uv\cos\theta+\sigma_{t}^{2}u^{2}\right)}.\end{split}$ (24)
Assume that a biological experiment regarding the response properties of the receptive field is performed by varying the angular frequency $\omega$ to get the maximum value of the response over this parameter. Differentiating the amplitude $A_{\varphi\varphi}$ with respect to $\omega$ and setting this derivative to zero, while also setting $\sigma_{2}=\kappa\,\sigma_{1}$ and $v=\sigma_{1}/\sigma_{t}$, then gives
$\hat{\omega}_{\varphi\varphi}=\frac{\sqrt{2}}{\sqrt{\frac{\sigma_{1}^{2}\left(\kappa^{2}v^{2}\sin^{2}(\theta)-2uv\cos(\theta)+2v^{2}\cos^{2}(\theta)+u^{2}\right)}{v^{2}}}}.$ (25)
Inserting this value into $A_{\varphi\varphi}(\theta,u,\omega;\;\sigma_{1},\sigma_{2},\sigma_{t},v)$, then gives the following combined direction and speed selectivity measure
$A_{\varphi\varphi,\max}(\theta,u,\;\kappa,v)=\\ =\frac{2v^{2}\cos^{2}(\theta)}{e\left(\kappa^{2}v^{2}\sin^{2}(\theta)-2uv\cos(\theta)+2v^{2}\cos^{2}(\theta)+u^{2}\right)},$ (26)
which for the inclination angle $\theta=0$ and for a coupling of the speed values of the form u=rvu=r\,v, reduces to the following speed dependency, notably again independent of the degree of elongation $\kappa$:
$R_{\varphi\varphi}(r)=A_{\varphi\varphi,\max}(0,r\,v;\;\kappa,r\,v)=\frac{1}{r^{2}-2r+2}.$ (27)
The middle left column in Figure 4 shows the result of plotting the measure $A_{\varphi\varphi,\max}(\theta,u;\;\kappa,v)$ of the direction selectivity as function of the inclination angle $\theta$ for a few values of the scale parameter ratio $\kappa$ for the special case when the speed of the stimulus is equal to the speed of the velocity $u=v$, with the values rescaled such that the peak value for each graph is equal to 1. As we can see from the graphs, the direction selectivity becomes sharper with increasing values of the degree of elongation $\kappa$. The middle left column in Figure 5 shows corresponding results of varying the magnitude of the motion according to u=rvu=r\,v for different values of $r$, while keeping the degree of elongation fixed to $\kappa=2$. As we can see from the results, again the direction selectivity of the simple cells prefers motion speeds $u$ that are near the preferred speed $v$ of the receptive field. The middle left graph in Figure 6 finally shows the relative speed sensitivity curve $R_{\varphi\varphi}(r)$.
4.3 Analysis for third-order simple cell without temporal differentiation
Let us next consider a velocity-adapted receptive field corresponding to a third-order scale-normalized Gaussian derivative with scale parameter $\sigma_{1}$ and velocity $v$ in the horizontal $x_{1}$-direction, a zero-order Gaussian kernel with scale parameter $\sigma_{2}$ in the vertical $x_{2}$-direction, and a zero-order Gaussian derivative with scale parameter $\sigma_{t}$ in the temporal direction, corresponding to $\varphi=0$, $v=0$, $\Sigma_{0}=\operatorname{diag}(\sigma_{1}^{2},\sigma_{2}^{2})$, $m=3$ and $n=0$ in (3.1):
$\begin{split}&T_{000,\scriptsize\mbox{norm}}(x_{1},x_{2},t;\;\sigma_{1},\sigma_{2},\sigma_{t})=\end{split}$
$\begin{split}&=\frac{\sigma_{1}^{3}}{(2\pi)^{3/2}\,\sigma_{1}\sigma_{2}\sigma_{t}}\,\partial_{x_{1}x_{1}x_{1}}\left.\left(e^{-x_{1}^{2}/2\sigma_{1}^{2}-x_{2}^{2}/2\sigma_{2}^{2}-t^{2}/2\sigma_{t}^{2}}\right)\right|_{x_{1}\rightarrow x_{1}-vt}\end{split}$
$\begin{split}&=\frac{-((x_{1}-vt)^{3}-3\sigma_{1}^{2}x)}{(2\pi)^{3/2}\,\sigma_{1}^{4}\sigma_{2}\sigma_{t}}\,e^{-(x_{1}-vt)^{2}/2\sigma_{1}^{2}-x_{2}^{2}/2\sigma_{2}^{2}-t^{2}/2\sigma_{t}^{2}}.\end{split}$ (28)
The corresponding receptive field response can then be expressed as, after solving the convolution integral in Mathematica,
$\begin{split}L_{000,\scriptsize\mbox{norm}}(x_{1},x_{2};\;\sigma_{1},\sigma_{2})=\end{split}$
$\begin{split}&=\int_{\xi_{1}=-\infty}^{\infty}\int_{\xi_{2}=-\infty}^{\infty}T_{000,\scriptsize\mbox{norm}}(\xi_{1},\xi_{2};\;\sigma_{1},\sigma_{2})\end{split}$
[…]
$\begin{split}&=-\omega^{3}\,\sigma_{1}^{3}\cos^{3}(\theta)\ \end{split}$
$\begin{split}&\phantom{==}\times e^{-\frac{\omega^{2}}{2}\left(\left(\sigma_{1}^{2}+\sigma_{t}^{2}v^{2}\right)\cos^{2}\theta+\sigma_{2}^{2}\sin^{2}\theta-2\sigma_{t}^{2}uv\cos\theta+\sigma_{t}^{2}u^{2}\right)}\end{split}$
$\begin{split}&\phantom{==}\times\cos(\cos(\theta)\,x_{1}+\sin(\theta)\,x_{2}-\omega\,u\,t+\beta),\end{split}$ (29)
i.e., it corresponds to sine wave with amplitude
$A_{\varphi\varphi\varphi}(\theta,u,\omega;\;\sigma_{1},\sigma_{2},v)=\omega^{3}\,\sigma_{1}^{3}\cos^{3}(\theta)\,\times\\ \times e^{-\frac{\omega^{2}}{2}\left(\left(\sigma_{1}^{2}+\sigma_{t}^{2}v^{2}\right)\cos^{2}\theta+\sigma_{2}^{2}\sin^{2}\theta-2\sigma_{t}^{2}uv\cos\theta+\sigma_{t}^{2}u^{2}\right)}$ (30)
To determine the value $\hat{\omega}$ of $\omega$ that gives the strongest response, let us differentiate $A_{\varphi}(\theta,u,\omega;\;\sigma_{1},\sigma_{2},v)$ with respect to $\omega$, set the derivative to zero, while also setting $\sigma_{2}=\kappa\,\sigma_{1}$ and $v=\sigma_{1}/\sigma_{t}$, which gives:
$\hat{\omega}_{\varphi\varphi\varphi}=\frac{\sqrt{3}}{\sqrt{\frac{\sigma_{1}^{2}\left(\kappa^{2}v^{2}\sin^{2}(\theta)-2uv\cos(\theta)+2v^{2}\cos^{2}(\theta)+u^{2}\right)}{v^{2}}}}.$ (31)
Inserting this value into $A_{\varphi\varphi\varphi}(\theta,u,\omega;\;\sigma_{1},\sigma_{2},v)$ then gives rise to a combined direction and speed selectivity measure of the form
$A_{\varphi\varphi\varphi,\max}(\theta,u;\;\kappa,v)=\\ =\frac{3\sqrt{3}\cos^{3}(\theta)}{e^{3/2}\left(\kappa^{2}\sin^{2}(\theta)-\frac{2u\cos(\theta)}{v}+2\cos^{2}(\theta)+\frac{u^{2}}{v^{2}}\right)^{3/2}},$ (32)
which for the inclination angle $\theta=0$ and for a coupling of the speed values of the form u=rvu=r\,v, reduces to the following speed dependency, notably again independent of the degree of elongation $\kappa$:
$R_{\varphi\varphi\varphi}(r)=A_{\varphi\varphi\varphi,\max}(0,r\,v;\;\kappa,r\,v)=\frac{1}{\left(r^{2}-2r+2\right)^{3/2}}.$ (33)
The middle right column in Figure 4 shows the result of plotting the measure $A_{\varphi\varphi\varphi,\max}(\theta,u;\;\kappa,v)$ of the direction selectivity as function of the inclination angle $\theta$ for a few values of the scale parameter ratio $\kappa$ for the special case when the speed of the stimulus is equal to the speed of the velocity $u=v$, with the values rescaled such that the peak value for each graph is equal to 1. As we can see from the graphs, the direction selectivity becomes sharper with increasing values of the degree of elongation $\kappa$. The middle right column in Figure 5 shows corresponding results of varying the magnitude of the motion according to u=rvu=r\,v for different values of $r$, while keeping the degree of elongation fixed to $\kappa=2$. As we can see from the results, again the direction selectivity of the simple cells prefers motion speeds $u$ that are near the preferred speed $v$ of the receptive field. The middle right graph in Figure 6 finally shows the relative speed sensitivity curve $R_{\varphi\varphi\varphi}(r)$.
4.4 Analysis for fourth-order simple cell without temporal differentiation
Let us next consider a fourth-order idealized model of a simple cell with with scale parameter $\sigma_{1}$ and velocity $v$ in the horizontal $x_{1}$-direction, a zero-order Gaussian kernel with scale parameter $\sigma_{2}$ in the vertical $x_{2}$-direction, and a zero-order Gaussian derivative with scale parameter $\sigma_{t}$ in the temporal direction, corresponding to $\varphi=0$, $v=0$, $\Sigma_{0}=\operatorname{diag}(\sigma_{1}^{2},\sigma_{2}^{2})$, $m=4$ and $n=0$ in (3.1):
$\begin{split}&T_{0000,\scriptsize\mbox{norm}}(x_{1},x_{2},t;\;\sigma_{1},\sigma_{2},\sigma_{t})=\end{split}$
$\begin{split}&=\frac{\sigma_{4}^{3}}{(2\pi)^{3/2}\,\sigma_{1}\sigma_{2}\sigma_{t}}\,\partial_{x_{1}x_{1}x_{1}x_{1}}\left.\left(e^{-x_{1}^{2}/2\sigma_{1}^{2}-x_{2}^{2}/2\sigma_{2}^{2}-t^{2}/2\sigma_{t}^{2}}\right)\right|_{x_{1}\rightarrow x_{1}-vt}\end{split}$
$\begin{split}&=\frac{-((x_{1}-vt)^{4}-6\sigma_{1}^{2}(x_{1}-vt)^{2}+3\sigma_{1}^{4})}{(2\pi)^{3/2}\,\sigma_{1}^{4}\sigma_{2}\sigma_{t}}\end{split}$
$\begin{split}&\phantom{=}\quad\times e^{-(x_{1}-vt)^{2}/2\sigma_{1}^{2}-x_{2}^{2}/2\sigma_{2}^{2}-t^{2}/2\sigma_{t}^{2}}.\end{split}$ (34)
After solving the convolution integral in Mathematica, the corresponding receptive field response is then of the form
$\begin{split}L_{0000,\scriptsize\mbox{norm}}(x_{1},x_{2};\;\sigma_{1},\sigma_{2})=\end{split}$
$\begin{split}&=\int_{\xi_{1}=-\infty}^{\infty}\int_{\xi_{2}=-\infty}^{\infty}T_{0000,\scriptsize\mbox{norm}}(\xi_{1},\xi_{2};\;\sigma_{1},\sigma_{2})\end{split}$
[…]
$\begin{split}&=\omega^{4}\,\sigma_{1}^{4}\cos^{4}(\theta)\ \end{split}$
$\begin{split}&\phantom{==}\times e^{-\frac{\omega^{2}}{2}\left(\left(\sigma_{1}^{2}+\sigma_{t}^{2}v^{2}\right)\cos^{2}\theta+\sigma_{2}^{2}\sin^{2}\theta-2\sigma_{t}^{2}uv\cos\theta+\sigma_{t}^{2}u^{2}\right)}\end{split}$
$\begin{split}&\phantom{==}\times\sin(\cos(\theta)\,x_{1}+\sin(\theta)\,x_{2}-\omega\,u\,t+\beta),\end{split}$ (35)
i.e., a sine wave with amplitude
$A_{\varphi\varphi\varphi\varphi}(\theta,u,\omega;\;\sigma_{1},\sigma_{2},v)=\\ =\omega^{4}\,\sigma_{1}^{4}\cos^{4}(\theta)\,e^{-\frac{1}{2}\omega^{2}(\sigma_{1}^{2}\cos^{2}\theta+\sigma_{2}^{2}\sin^{2}\theta)}.$ (36)
Again selecting the value of $\hat{\omega}$ at which the amplitude assumes its maximum over $\omega$ gives
$\hat{\omega}_{\varphi\varphi\varphi\varphi}=\frac{2}{\sqrt{\frac{\sigma_{1}^{2}\left(\kappa^{2}v^{2}\sin^{2}(\theta)-2uv\cos(\theta)+2v^{2}\cos^{2}(\theta)+u^{2}\right)}{v^{2}}}},$ (37)
which implies that the maximum amplitude over spatial scales as a function of the inclination angle $\theta$ and the scale parameter ratio $\kappa$ can be written
$A_{\varphi\varphi\varphi\varphi,\max}(\theta,u;\;\kappa,v)=\\ =\frac{16v^{4}\cos^{4}(\theta)}{e^{2}\left(\kappa^{2}v^{2}\sin^{2}(\theta)-2uv\cos(\theta)+2v^{2}\cos^{2}(\theta)+u^{2}\right)^{2}},$ (38)
which for the inclination angle $\theta=0$ and for a coupling of the velocity values of the form u=rvu=r\,v, reduces to the following speed dependency, notably again independent of the degree of elongation $\kappa$:
$R_{\varphi\varphi\varphi\varphi}(r)=A_{\varphi\varphi\varphi,\max}(0,r\,v;\;\kappa,v)=\frac{1}{\left(r^{2}-2r+2\right)^{2}}.$ (39)
The right column in Figure 4 shows the result of plotting the measure $A_{\varphi\varphi\varphi\varphi,\max}(\theta,u;\;\kappa,v)$ of the direction selectivity as function of the inclination angle $\theta$ for a few values of the scale parameter ratio $\kappa$ for the special case when the speed of the stimulus is equal to the speed of the velocity $u=v$, with the values rescaled such that the peak value for each graph is equal to 1. As we can see from the graphs, the direction selectivity becomes sharper with increasing values of the degree of elongation $\kappa$. The right column in Figure 5 shows corresponding results of varying the magnitude of the motion according to u=rvu=r\,v for different values of $r$, while keeping the degree of elongation fixed to $\kappa=2$. As we can see from the results, again the direction selectivity of the simple cells prefers motion speeds $u$ that are near the preferred speed $v$ of the receptive field. The right graph in Figure 6 finally shows the relative speed sensitivity curve $R_{\varphi\varphi\varphi\varphi}(r)$.
5 Direction and speed selectivity properties for idealized models of complex cells
Given the above analysis for a set of different models of simple cells, as varying by the order of spatial differentiation, let us next combine the output from these simple cells into a set of idealized models of complex cells.
$R_{{\cal Q}_{12}}(r)=\frac{e^{\frac{(r-1)^{2}-2\gamma^{2}}{\sqrt{2}\left(r^{2}-2r+2\right)}}\sqrt{\left(r^{2}-2r+3\right)e^{\frac{2\sqrt{2}\gamma^{2}}{r^{2}-2r+2}}+(r-1)^{2}}}{\sqrt{2}\left(r^{2}-2r+2\right)}$ (40)
$R_{{\cal Q}_{34}}(r)=\sqrt{\frac{e^{\frac{2\sqrt{3}\left(2\gamma^{2}+1\right)(r-1)^{2}}{r^{2}-2r+2}}\left(\left(r^{2}-2r+\sqrt{6}+2\right)e^{\frac{4\sqrt{3}\gamma^{2}}{r^{2}-2r+2}}+r^{2}-2r-\sqrt{6}+2\right)}{\left(\left(1+\sqrt{6}\right)e^{4\sqrt{3}\gamma^{2}}+1-\sqrt{6}\right)\left(r^{2}-2r+2\right)^{4}}}$ (41)
Figure 7: Dependencies on the relative speed parameter $r$ for the integrated quasi quadrature measures $\overline{{\cal Q}}_{\varphi,12,\scriptsize\mbox{vel},\scriptsize\mbox{norm}}L$ according to (12) and $\overline{{\cal Q}}_{\varphi,34,\scriptsize\mbox{vel},\scriptsize\mbox{norm}}L$ according to (13), when subjected to a moving sine wave with inclination angle $\theta=0$. The parameter $\gamma$ is the relative integration scale factor for the spatial integration of the corresponding pointwise quasi quadrature measure with an affine Gaussian kernel.
Velocity-tuned complex cell (12) Velocity-tuned complex cell (34)
$\kappa=1$ Refer to caption Refer to caption
$\kappa=2$ Refer to caption Refer to caption
$\kappa=4$ Refer to caption Refer to caption
$\kappa=8$ Refer to caption Refer to caption
Figure 8: Graphs of the direction selectivity for idealized models of complex cells based on (left column) spatial integration of a quadratic combination of first- and second-order spatial derivatives with zero-order temporal Gaussian kernels into $\overline{{\cal Q}}_{\varphi,12,\scriptsize\mbox{vel},\scriptsize\mbox{norm}}L$ according to (12) (right column) spatial integration of a quadratic combination of third- and fourth-order spatial derivatives with zero-order temporal Gaussian kernels into $\overline{{\cal Q}}_{\varphi,34,\scriptsize\mbox{vel},\scriptsize\mbox{norm}}L$ according to (13), shown for different values of the degree of elongation $\kappa$ between the spatial scale parameters in the vertical vs. the horizontal directions. Observe how the direction selectivity varies strongly depending on the eccentricity $\epsilon=1/\kappa$ of the receptive fields. (top row) Results for $\kappa=1$. (second row) Results for $\kappa=2$. (third row) Results for $\kappa=4$. (bottom row) Results for $\kappa=8$.
Velocity-tuned complex cell (12) Velocity-tuned complex cell (34)
Refer to caption Refer to caption
Figure 9: Graphs of the speed selectivity curves $R_{{\cal Q}_{12}}(r)$ according to (40) and $R_{{\cal Q}_{34}}(r)$ according to (41) for the inclination angle $\theta=0$ for velocity-tuned models of complex cells cells based on (left) an integrated quadratic combination of first- and second-order directional derivatives of affine Gaussian kernels combined with zero-order temporal Gaussian kernels, (right) an integrated quadratic combination of third- and fourth-order directional derivatives of affine Gaussian kernels combined with zero-order temporal Gaussian kernels, shown for different values of the parameter $r$ in the relationship u=rvu=r\,v between the speed $u$ of the motion stimulus and the speed $v$ of the receptive field. Notably, these measures are independent of the degree of elongation $\kappa$ of the spatio-temporal receptive fields. (Horizontal axes: relative velocity parameter $r\in[0,4]$. Vertical axes: velocity sensitivity function $R_{\cal Q}(r)$.)
5.1 Analysis for integrated velocity-tuned models of complex cells
To model the spatial response of a complex cell according to the spatio-temporal quasi-quadrature measure (6) based on velocity-adapted spatio-temporal receptive fields, we combine either the responses of the first- and second-order simple cells
$({\cal Q}_{0,12,\scriptsize\mbox{vel},\scriptsize\mbox{norm}}L)^{2}=L_{0,\scriptsize\mbox{norm}}^{2}+\,C_{\varphi}\,L_{00,\scriptsize\mbox{norm}}^{2}$ (42)
with $L_{0,\scriptsize\mbox{norm}}$ according to (LABEL:eq-L0-vel-adapt-anal) and $L_{00,\scriptsize\mbox{norm}}$ according to (LABEL:eq-L00-vel-adapt-anal). Similarly, we model the spatial response of a complex cell according to the spatio-temporal quasi-quadrature measure (7) based on the responses of the third- and fourth-order simple cells as
$({\cal Q}_{0,34,\scriptsize\mbox{vel},\scriptsize\mbox{norm}}L)^{2}=L_{000,\scriptsize\mbox{norm}}^{2}+\,C_{\varphi\varphi\varphi}\,L_{0000,\scriptsize\mbox{norm}}^{2},$ (43)
with $L_{000,\scriptsize\mbox{norm}}$ according to (LABEL:eq-L000-vel-adapt-anal) and $L_{0000,\scriptsize\mbox{norm}}$ according to (LABEL:eq-L0000-vel-adapt-anal).
Selecting the angular frequency as the geometric average of the angular frequency values at which the above spatio-temporal simple cell models assume their maxima over angular frequencies, gives either
$\hat{\omega}_{{\cal Q}_{12}}=\sqrt{\hat{\omega}_{\varphi}\,\hat{\omega}_{\varphi\varphi}}\\ =\sqrt[4]{2}\times\\ \sqrt{\frac{v^{2}}{\sigma_{1}^{2}\left(\kappa^{2}v^{2}\sin^{2}(\theta)-2uv\cos(\theta)+2v^{2}\cos^{2}(\theta)+u^{2}\right)}}$ (44)
with $\hat{\omega}_{\varphi}$ according to (19) and $\hat{\omega}_{\varphi\varphi}$ according to (25), or
$\hat{\omega}_{{\cal Q}_{34}}=\sqrt{\hat{\omega}_{\varphi\varphi\varphi}\,\hat{\omega}_{\varphi\varphi\varphi\varphi}}=\\ =\sqrt{2}\sqrt[4]{3}\times\\ \sqrt{\frac{v^{2}}{\sigma_{1}^{2}\left(\kappa^{2}v^{2}\sin^{2}(\theta)-2uv\cos(\theta)+2v^{2}\cos^{2}(\theta)+u^{2}\right)}}$ (45)
with $\hat{\omega}_{\varphi\varphi\varphi}$ according to (31) and $\hat{\omega}_{\varphi\varphi\varphi\varphi}$ according to (37).
Integrating the resulting expressions in Mathematica, with an affine Gaussian window function according to (12) or (13), and setting the relative weights between first- and second-order information or between third- and fourth-order information to $C_{\varphi}=1/\sqrt{2}$ and $C_{\varphi\varphi\varphi}=1/\sqrt{2}$, while also setting $\sigma_{2}=\kappa\,\sigma_{1}$ and $v=\sigma_{1}/\sigma_{t}$, then leads to the resulting composed amplitude measures $A_{{\cal Q}_{12},\max}(\theta,u;\;\kappa,v)$ and $A_{{\cal Q}_{34},\max}(\theta,u;\;\kappa,v)$ as functions of the inclination angle $\theta$, the speed $u$ of the input stimulus, the degree of elongation $\kappa$ and the velocity parameter $v$ in closed form. Unfortunately, these expressions are, however, too complex to be reproduced here.
In the special case when the inclination angle $\theta=0$, and we parameterize the speed parameters $u$ and $v$ according to u=rvu=r\,v, the resulting dependencies on the parameter $r$ can, however, be reproduced, as shown in Figure 7.
Figure 8 shows examples of orientation selectivity curves that arise from the explicit expressions that we have derived for the composed amplitude measures $A_{{\cal Q}_{12},\max}(\theta,u;\;\kappa,v)$ and $A_{{\cal Q}_{34},\max}(\theta,u;\;\kappa,v)$, in the special case when the speed parameter $u$ of the motion stimulus is equal to the speed parameter $v$ of the spatio-temporal receptive field in the horizontal direction. As can be seen from these results, the direction selectivity properties do, as for the previous analysis of simple cells, become significantly sharper with increasing orders of spatial differentiation as well as with increasing degree of elongation $\kappa$ of the receptive fields.
Figure 9 shows corresponding graphs of the speed sensitivity curves $R_{{\cal Q}_{12}}(r)$ according to (40) and $R_{{\cal Q}_{34}}(r)$ according to (41) for the inclination angle $\theta=0$. As can be seen from these results, the speed sensitivity decreases more abruptly towards larger relative speed factors $r$ with increasing order of spatial differentiation in the idealized models of the spatio-temporal receptive fields.
$A_{\varphi}(\theta,u;\;\sigma_{1},\kappa,\sigma_{t},v)=\frac{\cos(\theta)\sqrt{\sigma_{1}^{2}+\sigma_{t}^{2}(u-v)^{2}}}{\sqrt{\kappa^{2}\sigma_{1}^{2}\sin^{2}(\theta)+\cos^{2}(\theta)\left(\sigma_{1}^{2}+\sigma_{t}^{2}v^{2}\right)-2\sigma_{t}^{2}uv\cos(\theta)+\sigma_{t}^{2}u^{2}}}$ (46)
$A_{\varphi\varphi}(\theta,u;\;\sigma_{1},\kappa,\sigma_{t},v)=\frac{\cos^{2}(\theta)\left(\sigma_{1}^{2}+\sigma_{t}^{2}(u-v)^{2}\right)}{\kappa^{2}\sigma_{1}^{2}\sin^{2}(\theta)+\cos^{2}(\theta)\left(\sigma_{1}^{2}+\sigma_{t}^{2}v^{2}\right)-2\sigma_{t}^{2}uv\cos(\theta)+\sigma_{t}^{2}u^{2}}$ (47)
$A_{\varphi\varphi\varphi}(\theta,u;\;\sigma_{1},\kappa,\sigma_{t},v)=\frac{\cos^{3}(\theta)\left(\sigma_{1}^{2}+\sigma_{t}^{2}(u-v)^{2}\right)^{3/2}}{\left(\kappa^{2}\sigma_{1}^{2}\sin^{2}(\theta)+\cos^{2}(\theta)\left(\sigma_{1}^{2}+\sigma_{t}^{2}v^{2}\right)-2\sigma_{t}^{2}uv\cos(\theta)+\sigma_{t}^{2}u^{2}\right)^{3/2}}$ (48)
$A_{\varphi\varphi\varphi\varphi}(\theta,u;\;\sigma_{1},\kappa,\sigma_{t},v)=\frac{\cos^{4}(\theta)\left(\sigma_{1}^{2}+\sigma_{t}^{2}(u-v)^{2}\right)^{2}}{\left(\kappa^{2}\sigma_{1}^{2}\sin^{2}(\theta)+\cos^{2}(\theta)\left(\sigma_{1}^{2}+\sigma_{t}^{2}v^{2}\right)-2\sigma_{t}^{2}uv\cos(\theta)+\sigma_{t}^{2}u^{2}\right)^{2}}$ (49)
Figure 10: Explicit expressions for the direction selectivity curves for idealized models of simple cells for different orders $m\in\{1,2,3,4\}$ of spatial differentiation, without any coupling between the spatial scale parameter $\sigma_{1}$ and the temporal scale parameter $\sigma_{t}$.
6 Condensed summary of direction and speed selectivity properties for the idealized models of simple cells and complex cells
By summarizing results for the direction selectivity of the different idealized models of simple cells in Section 4, we find that provided that the spatial and the temporal scale parameters $\sigma_{1}$ and $\sigma_{t}$ are coupled to the velocity parameter $v$ of the spatio-temporal receptive field according to $\sigma_{1}/\sigma_{t}=v$, the combined direction and speed selectivity is of the form
$A_{\varphi^{m},\max}(\theta,u;\;\kappa,v)=\\ =\alpha_{m}\left(\frac{\cos(\theta)}{\sqrt{\kappa^{2}\sin^{2}(\theta)-\frac{2u\cos(\theta)}{v}+2\cos^{2}(\theta)+\frac{u^{2}}{v^{2}}}}\right)^{m},$ (50)
as depending on the order $m$ of spatial differentiation in the spatial component of the idealized receptive field model. Specifically, in the special case when the inclination angle $\theta=0$, the resulting speed selectivity is for u=rvu=r\,v of the form
$R_{\varphi^{m}}(r)=A_{\varphi^{m},\max}(0,r\,v;\;\kappa,v)=\frac{1}{\left(r^{2}-2r+2\right)^{m/2}}.$ (51)
Corresponding combined direction and speed selectivity measures for the more general case, when the spatial and the temporal scale parameters $\sigma_{1}$ and $\sigma_{t}$ are not coupled in relation to the speed parameter $v$ of the receptive field, are given in Equations (46)–(49) in Figure 10. In the special case when the inclination angle $\theta=0$, these expressions reduce to resulting speed selectivity curves of the form
$R_{\varphi^{m}}(u;\;\sigma_{1},\sigma_{t},v)=\left(\frac{\sigma_{1}}{\sqrt{\sigma_{1}^{2}+\sigma_{t}^{2}\,(u-v)^{2}}}\right)^{m}.$ (52)
Regarding qualitative properties of the direction selectivity for the different orders $m$ of spatial differentiation, by comparing the direction selectivity graphs for the idealized models of simple cells, we generally find that:
  • The direction selectivity becomes sharper with increasing degree of elongation $\kappa$ and for increasing order $m$ of spatial differentiation.
  • The decrease in speed sensitivity with increasing speed of the probing motion stimulus becomes more abrupt with increasing order of spatial differentiation.
While the expressions for idealized models of complex cells are too complex to be reproduced in this treatment, corresponding qualitative properties do also hold regarding the idealized models of velocity-tuned complex cells analyzed in Section 5.
7 Relations to direction and speed selectivity properties of biological neurons
In this section, we will relate the above derived theoretical relationships of the direction selectivity properties of spatio-temporal receptive fields according to the generalized Gaussian derivative model for visual receptive fields to available neurophysiological recordings of motion-sensitive neurons in the primary visual cortex.
7.1 Relations to neurophysiological measurements of motion-sensitive neurons in the primary visual cortex
In their in-depth analysis of velocity sensitivity and direction selectivity properties of neurons in the primary visual cortex (V1) in monkeys, Orban et al. (1986) found that [^1] The terminology “speed” that we use in this paper means the same physical entity as Orban et al. (1986) refer to as “velocity”. The reason why we use speed in this paper, is that according to the convention in physics, the notion of “speed” is without any direction, while the term “velocity” comprises both the speed and the direction of motion.
  • “Velocity sensitivity of V1 neurons shifts to faster velocities with increasing eccentricity.”
  • “All neurons could be classified into three categories according to their velocity-response curves: velocity low pass, velocity broad band, and velocity tuned.”
  • “There is a significant correlation between velocity upper cutoff and receptive field width among V1 neurons. The change in upper cutoff velocity with eccentricity depends both on temporal and spatial factors.”
Since our way of probing the idealized models of simple cells and complex cells by sine waves differs from the way that above neurophysiological results have been obtained using moving bars, we cannot expect to be able to directly compare [^2] To perform more quantitative comparisons to the neurophysiological results, it would furthermore be necessary to get access to raw data for a larger population of visual neurons, which we do not have any access to. our direction and speed selectivity curves to the direction and speed selectivity curves recorded by Orban et al. (1986). We can, however, note that the idealized models of simple cells in Section 3.1 and that our idealized models of complex cells in Section 3.2, both based on velocity-adapted Gaussian derivative operators, have qualitative similarities to the class of velocity-tuned neurons in V1 characterized by Orban et al. (1986), in the sense of being selective to specific directions and speeds of the input stimulus.
By choosing the speeds of the motion parameters in the idealized models of simple cells and complex cells sufficiently low, their resulting speed selectivity properties would be structurally similar to the motion sensitive neurons with speed low pass character, as described by Orban et al. (1986). Furthermore, by combining the output from such velocity-tuned models of complex cells over wider ranges of motion speeds, while keeping the direction of motion the same, it should also be possible to formulate idealized models of complex cells with more broadband speed selectivity properties, as described by Orban et al. (1986). In these ways, there are therefore qualitative similarities between the direction and speed selectivity properties of our idealized models of simple cells and the neurophysiological results by Orban et al. (1986), thus supporting the idea of expanding theoretical models of motion-sensitive neurons with respect to the degrees of freedom of local Galilean transformations.
7.2 Relationships to the hypothesis about Galilean covariant spatio-temporal receptive fields in the primary visual cortex
In relation to covariance properties of visual receptive fields with regard to basic geometric image transformations, Lindeberg (2023b; 2025d) proposed the working hypothesis that simple cells in primary visual cortex ought to be covariant with respect to Galilean transformations of the form
$\left\{\begin{array}[]{l}x_{1}^{\prime}=x_{1}+v_{1}\,t,\\ x_{2}^{\prime}=x_{2}+v_{2}\,t.\end{array}\right$. (53)
This would then imply that the shapes of the spatio-temporal receptive fields of simple cells in the primary visual cortex ought to be expanded over the degrees of freedom of Galilean transformations. Concretely, this would imply that spatio-temporal receptive fields of the form (3.1) would be present for a range of image velocities $v=(v_{1},v_{2})^{T}\in{\mathbb{R}}^{2}$. One may therefore ask if there would be neurophysiological support for this hypothesis.
First of all, the comparison between our theoretical results regarding direction and speed selectivity properties of our idealized models of simple cells and complex cells to the neurophysiological results by Orban et al. (1986) in Section 7.1 are consistent with an expansion of the spatio-temporal receptive field shapes over the degrees of freedom of local Galilean transformations.
Furthermore, regarding the processing of time-dependent image data in the primary visual cortex and the middle temporal visual area MT, we have the following facts:
  • the ability of the simple cells in the visual system of higher mammals to compute spatio-temporal receptive field responses similar to velocity-adapted temporal derivatives (DeAngelis et al. 1995; 2004; Lindeberg 2021 Figure 18 bottom part),
  • visual neurons in the primate visual cortex being known to be direction selective (Hubel 1959, Orban et al. 1986, Churchland et al. 2005), and “… organized into subcolumns within an iso-orientation column, with each subcolumn preferring motion in a different direction” (Elstrott and Feller 2009),
  • the receptive fields in the middle temporal visual area MT being able to compute to a rich variety of direction-selective responses (Orban 1997, Born and Bradley 2005),
  • with the output from the primary visual cortex providing input to the middle temporal visual area MT (Movshon and Newsome 1996) and
  • with the direction-selective neurons in MT organized into direction columns (Albright et al. 1984).
From these results, it appears as if we can regard the primary visual cortex V1 and the middle temporal visual area MT as performing expansions of the visual representations over the directions and magnitudes of local Galilean motions. Thereby, it seems plausible that:
  • the visual system should be able to compute Galilean-covariant receptive field responses, so as to be able to process visual information over a wide range of motion speeds and motion directions,
  • this motion processing hierarchy could then be based on Galilean-covariant receptive fields in the primary visual cortex.
Given such an expansion of the shapes of the shapes of the spatio-temporal receptive fields in the primary visual cortex over the degrees of freedom of local Galilean transformations, a consequence of that would then be variabilities in the direction selectivity properties and the speed sensitivity properties of the visual neurons, as recorded in neurophysiological measurements.
8 Suggestions for further neurophysiological experiments
Given the above theoretical results, with their qualitative relationships to existing results from neurophysiological measurements, a highly interesting topic would be to aim at additional more quantitative comparisons.
The publicly available data regarding explicitly reconstructed spatio-temporal receptive fields from neurophysiological recordings is, however, very limited. For this reason, it would be highly interesting if complementary neurophysiological recordings [^3] Such experiments should preferably be done for higher mammals with general purpose vision systems, like primates or cats. Recordings from mice may specifically not be appropriate, since the vision system of mice is far less developed compared the visual systems of higher mammals, see e.g. Huberman and Niell (2011). Specifically, recent results by Fu et al. (2024) of reconstructing most exciting inputs (MEIs) for early visual receptive fields in mice suggest that complex spatial features emerge earlier in the visual pathway of mice compared to primates, and that the receptive fields of mice are therefore more complex than a Gaussian derivative model, as used in the theoretical model for performing the theoretical analysis in this paper. could be performed, if possible with recordings of both direction and speed selectivity properties of visual neurons with explicitly reconstructed spatio-temporal receptive fields from the same neurons, to answer the following theoretically motivated questions:
  • How well can the spatio-temporal receptive fields of simple cells in the primary visual cortex be modelled by velocity-adapted spatio-temporal affine-Gaussian-derivative-based receptive fields of the form (3.1), if the entire 2+1-D spatio-temporal receptive field is reconstructed, and not only visualized in a 1+1-D cross-section, as done by DeAngelis et al. (1995; 2004) and de Valois et al. (2000), and for which the match is qualitatively very good, see Lindeberg 2021 Figure 18 bottom part?
  • Which orders $m$ of spatial differentiation describe the variabilities in spatio-temporal receptive fields of simple cells in the primary visual cortex?
  • How wide is the variability in the degree of elongation $\kappa$ in such models of the spatio-temporal receptive fields of simple cells in the primary visual cortex?
  • How wide are the variabilities in the spatial scales $\sigma_{1}$ and the temporal scales $\sigma_{t}$, as depending on the distance from the center of the fovea? Does the vision system only implement a set of finest spatial scales $\sigma_{1}$, or does it comprise a substantial variability? Similarly, does the vision system only implement a set of finest temporal scales $\sigma_{t}$? How is the ratio $\sigma_{1}/\sigma_{t}$ related to the speed parameter $|v|$ of the local Galilean motion in the velocity-adapted model of the spatio-temporal receptive fields?
  • How well do the neurophysiological recordings of the direction and speed selectivities of the visual neurons agree with the theoretical predictions obtained from first determining the shape parameter $\varphi$, $\sigma_{1}$, $\kappa$, $\sigma_{t}$ and $v$ in an idealized model of the spatio-temporal receptive field, and then using the predictions about direction and speed selectivity properties from the here presented theoretical results?
Answering these questions would provide further cues towards understanding the computational function of the spatio-temporal receptive fields in the primary visual cortex, and concerning to what extent we can regard the influence of geometric image transformations on the receptive field responses as a primary factor for the development of the visual receptive fields. Such more detailed comparisons between the properties of the theoretical model and the properties of the biological neurons could also provide cues to possible ways of extending the theoretical model for spatio-temporal receptive fields with additional [^4] The current model for linear spatio-temporal receptive fields, that we base this work upon, has been derived under minimal assumptions regarding the structure of the environment in combination with internal consistency requirements to guarantee theoretically well-founded treatment of image structures over multiple spatial and temporal scales (Lindeberg 2011; 2021). a priori information, if motivated from results from neurophysiological experiments.
9 Summary and discussion
We have presented a theory for how the direction and speed selectivity properties of idealized models of simple cells and complex cells can be related to properties of the underlying spatio-temporal receptive fields in such models of motion-sensitive visual neurons.
These receptive field models have been obtained from a normative theory of visual receptive fields, derived from axiomatic assumptions regarding structural properties of the environment in combination with internal consistency requirements to guarantee theoretically well-founded processing of image structures over multiple spatial and temporal scales. This theory states that a canonical way to model linear receptive fields, corresponding to simple cells in the primary visual cortex, is in terms of spatial and temporal derivatives of affine Gaussian kernels, combined with a mechanism of velocity adaptation to handle the influence on image data from local image motions. Based on these idealized models of simple cells, we have formulated idealized models of complex cells, based on local quadratic energy models of simple cells, combined with regional spatial integration.
By subjecting such idealized models of motion-sensitive neurons in the primary visual cortex to a structurally similar probing method as used for probing the direction and speed selectivity properties of visual neurons, we have derived closed form expressions for the direction and speed selectivity properties of our idealized models of simple cells, as summarized in Equations (50) and (51) for the special case when spatial and temporal scale parameters $\sigma_{1}$ and $\sigma_{t}$ are coupled to the speed parameter $v$ of the spatio-temporal receptive field according to $\sigma_{1}/\sigma_{t}=v$. We have also derived more general results for such combined direction and speed selectivity properties, reproduced in Figure 10 and Equation (52) for the more general case, when there is no coupling between the scale parameters in relation to the speed parameter of the receptive field.
In essence, these results show that the direction selectivity properties become more narrow with increasing order of spatial differentiation as well as with increasing degree of elongation in the underlying spatial components of the receptive fields. The speed selectivity properties also become more narrow with increasing order of spatial differentiation, but are for the inclination angle $\theta=0$ independent of the degree of elongation of the receptive fields.
For our idealized models of complex cells, the resulting closed form expressions of the direction selectivity properties are unfortunately too complex to be reported here. The qualitative results regarding more narrow direction selectivity properties with increasing order of spatial integration and increasing degree of elongation do, however, still hold. The speed selectivity properties for our idealized models of complex cells for the inclination angle $\theta=0$ are reported in Figure 7.
By comparisons with the results of neurophysiological recordings of motion-sensitive neurons in the primary visual cortex in Section 7.1, the presented theoretical results are qualitatively consistent with the previously reported class of velocity-tuned motion sensitive neurons, in the sense that the models of the neurons respond maximally to the combination of a preferred motion direction and a preferred motion speed, and that the magnitude of the response decreases when changing the motion direction and/or the motion speed from the preferred value.
By comparisons with overall results concerning direction selective neurons in the primary visual cortex V1 and the middle temporal area MT in Section 7.2, with their organization into subcolumns preferring motion in different directions, we have also found this organization consistent with the hypothesis proposed in Lindeberg (2023b; 2025d), that the population of receptive fields of the simple cells in the primary visual cortex ought to be covariant under local Galilean transformations.
This would then specifically imply that the spatio-temporal receptive ought to have their shapes expanded over the degrees of freedom of local transformations. In practice, this would imply that there would be multiple copies of similar types of spatio-temporal receptive fields for different values of the velocity parameter $v$ in the idealized model (3.1) of simple cells, as illustrated in Figures 2 and 2.
Unfortunately, the amount of publicly available data regarding the neurophysiological properties of spatio-temporal receptive fields in the primary visual cortex is very limited. To address further questions regarding the quantitative modelling of the spatio-temporal receptive fields of simple cells and complex cells, with their direction and speed selectivity properties, we have therefore proposed a set of outlines for further neurophysiological measurements in Section 8.
Notwithstanding such potential opportunities for additional more detailed quantitative modelling and comparisons if further neurophysiological measurements would become available, we have in this treatment theoretically analyzed how the direction and speed selectivity properties of idealized models of simple cells and complex cells can be related to inherent properties of the their underlying spatio-temporal receptive fields, and derived explicit relationships for how the direction and speed selectivity properties can for these models be directly related to intrinsic shape parameters of the receptive fields. Our intention is that these results should contribute to a better theoretical understanding regarding the computational mechanisms underlying the processing of motion stimuli in the primary visual cortex.
Overall, these results are also consistent with the wider hypothesis in Lindeberg (2021; 2023b; 2025d) that the influence of geometric image transformations on the receptive field responses may be a primary factor in the development of the receptive fields in the primary visual cortex.